# Many-Task Federated Fine-Tuning via Unified Task Vectors

### v1

@@ Abstract
Federated Learning (FL) traditionally assumes homogeneous client tasks; however, in real-world scenarios, clients often specialize in diverse tasks, introducing task heterogeneity. To address this challenge, Many-Task FL (MaT-FL) has emerged, enabling clients to collaborate effectively despite task diversity. Existing MaT-FL approaches rely on client grouping or personalized layers, requiring the server to manage individual models and failing to account for clients handling multiple tasks. We propose MaTU, a MaT-FL approach that enables joint learning of task vectors across clients, eliminating the need for clustering or client-specific weight storage at the server. Our method introduces a novel aggregation mechanism that determines task similarity based on the direction of clients’ task vectors and constructs a “unified” task vector encapsulating all tasks. To address task-specific requirements, we augment the “unified” task vector with lightweight modulators that facilitate knowledge transfer among related tasks while disentangling dissimilar ones. Evaluated across 30 datasets, MaTU achieves superior performance over state-of-the-art MaT-FL approaches, with results comparable to per-task fine-tuning, while delivering significant communication savings.
@@ 1 Introduction
Recent studies Chen et al. (2023); Zhuang et al. (2023); Cai et al. (2023) have underscored the significance of task heterogeneity in FL. To address this diversity, traditional FL has evolved into Many-Task Federated Learning (MaT-FL) Cai et al. (2023) to enable clients to effectively collaborate, despite specializing in different tasks. It is worth to mention that while “many-task” has been interchangeably used as “multi-task” in conventional ML, early Multi-Task Federated Learning (MTFL) mainly addressed personalized FL, a subset of MaT-FL. MaT-FL is thus used to distinguish it from prior MTFL works Cai et al. (2023); Muhamed et al. (2024); Zhuang et al. (2023). Current MaT-FL approaches address task heterogeneity by focusing on client grouping or split FL. For instance, FedBone Chen et al. (2023) employs split FL whereas the server needs to sustain a unique model per client. Conversely, MaT-FL Cai et al. (2023) and MAS Zhuang et al. (2023) concentrate on dynamically grouping client models based on task similarity, aggregating models among clients that handle similar tasks. Nevertheless, these strategies necessitate the server managing individual models for each client or task, while they do not consider cases in which each client hold more than one task. Nevertheless, these strategies require the server to manage individual models for each client or task and fail to address scenarios where clients hold more than one task (e.g., a self-driving car handling tasks like object detection, lane recognition, and pedestrian tracking simultaneously). Training a single model among tasks offers considerable benefits in federated settings. Training individual models for each task is resource-intensive, whereas a unified model not only conserves resources but also potentially enhances performance by leveraging learning from diverse auxiliary tasks. This capability is particularly vital in pragmatic FL scenarios in which data availability dramatically varies across tasks — some enjoying ample data, while others are data-starved.
@@
This paper addresses the challenge of adapting pretrained Foundation Models (FMs) to the MaT-FL setting, where clients hold an arbitrary (>$1$) number of tasks, and we aim to train a unified model across all tasks. We propose Many-Task FL via Unified Task Vectors (MaTU), a novel approach that eliminates the need for clustering or client-specific weight storage at the server, enabling natural knowledge transfer across tasks in FL. Inspired by recent advancements in model merging and Task Arithmetic (TA) Ilharco et al. (2023); Yadav et al. (2023a)MaTU introduces a task vector aggregation mechanism tailored for MaT-FL. Specifically, we determine task similarity directly based on the direction of clients’ task vectors and constructs a “unified” task vector that encapsulates all tasks. To handle task-specific requirements, we augment the “unified” vector with lightweight modulators (i.e., binary task-specific masks and scalers) that facilitate knowledge transfer between related tasks while promoting weight disentanglement for dissimilar ones, ensuring robust performance across heterogeneous tasks in FL. Concisely, our main contributions are as follows:
@@ 2 Related Work
MaT-FL. Many-Task Federated Learning (MaT-FL) has emerged to tackle task heterogeneity in FL. Although the term “many-task” mirrors “multi-task” in conventional ML, early Multi-Task Federated Learning (MTFL) primarily focused on personalized FL, a specific case of MaT-FL. Therefore, we adopt the term MaT-FL to distinguish it from earlier MTFL works, as seen in Cai et al. (2023); Muhamed et al. (2024); Zhuang et al. (2023). From MTFL approaches, FedProx Li et al. (2020), introduced a proximity term to limit client updates from deviating too far from the global model in MTFL. Alternatively, FedBone Chen et al. (2023) used split FL to introduce task-specific personalized layers to handle multiple tasks in FL.
Recent MaT-FL approaches primarily address task heterogeneity by focusing on client grouping Cai et al. (2023); Zhuang et al. (2023); Lu et al. (2023). MaT-FL Cai et al. (2023) and MAS Zhuang et al. (2023) dynamically group clients based on task similarity, enabling model aggregation among clients with similar tasks. FedHCA Lu et al. (2023) extends this by supporting a variable number of tasks across clients, creating a more flexible framework. More recently, NTKFedAvg Muhamed et al. (2024) introduced task arithmetic in MaT-FL, enabling clients to train task-specific adapters and leverage server-side adapter fusion for optimizing multiple tasks. To further improve task disentanglement during model aggregation, they applied Neural Tangent Kernel (NTK) linearization over the model prior to training. Nevertheless, group-based MaT-FL approaches require the server to manage multiple models for each task group, assume task relationships are known in advance, and introduce significant client-side overhead, making scalability a challenge as the number of tasks increases. Training a single model across tasks offers considerable benefits: it reduces resource consumption and can enhance performance by leveraging diverse auxiliary tasks. While NTKFedAvg attempts to address these challenges, it focuses on a single task per client, whereas in practice, clients may hold multiple tasks. MaTU bridges this gap by training a unified model across all clients and tasks, dynamically building task correlations to enable knowledge transfer, and using task vector aggregation based on task similarity, removing the need for multiple server-side models.
@@ 3.1 Preliminaries
For the set of tasks assigned to each clients, $\mathcal{T}_{n}$, we can now utilize the newly computed aggregated task-specific vector to derive their updated “unified” task vector, $\tau^{r+1}_{n}$, and subsequently compute the set of task-specific masks, $\mathcal{M}^{r+1}_{n}=\{(\tau^{i,r+1}\odot\tau^{r+1}_{n}>0)\}^{k_{n}}_{i=1}$, and scaling parameters $\lambda^{r+1}_{n}=\left\{\frac{\sum|\tau^{i,r+1}_{n}|}{\sum|m^{i,r+1}_{n}\odot\tau^{r+1}_{n}|}\right\}_{i=1}^{k_{n}}$. The client’s “unified” task vector, $\tau^{r+1}_{n}$, and task-specific modules, $\mathcal{M}^{r+1}_{n}$ and $\lambda^{r+1}_{n}$, are then transmitted to the client-side for subsequent training. To train task $t$, the client must modulate the “unified” vector as $\dot{\tau}^{r+1}_{t}=\lambda^{t,r+1}_{n}\cdot m^{t,r+1}_{n}\odot\tau^{r+1}_{n}$, and injects it to the pre-trained model (i.e., $\theta^{r+1}_{t}=\theta_{p}+\dot{\tau}^{r+1}_{t}$). Note that, while it can be done on the server-side, performing the modulation locally on clients results in significant communication savings and is computationally cheap.
@@ 4.1 Results
$8$-task benchmark: Table 2 demonstrates that MaTU maintains strong performance in multiple-task settings, achieving an average accuracy of 79.47% – a modest 5% drop from single-task settings – while outperforming all baselines and effectively minimizing task interference via masking. Traditional FL methods, such as FedAvg and FedProx, exhibit significant performance deficits, with average accuracies well below their single-task counterparts, underscoring their limitations in addressing task heterogeneity. Notably, NTK-FedAvg surpasses these baselines, likely benefiting from improved task disentanglement through NTK-based linearization. In stark contrast, FedPer experiences a severe performance decline ($\approx$40%) when transitioning from single to multiple-task settings, reflecting its inability to manage intra-client task heterogeneity due to its reliance on personalization layers. This highlights that personalized FL approaches are not designed to address the broader challenges of task heterogeneity, underscoring that personalization represents a subset of the more general MaT-FL problem. Conversely, MaT-FL achieves higher performance by employing similarity-based grouping to mitigate task conflicts, reducing interactions among dissimilar clients and tasks, but still lags MaTU by 10%. We further analyze how MaT-FL and MaTU handle such conflicts in Fig.˜6.
Beyond superiority in terms of test set model accuracy, MaTU demonstrates exceptional communication efficiency in multiple-task settings, as illustrated by the $\mathrm{bpt}$ column in Table 2. Competing methods transmit multiple adapters — one for each task held by a client — resulting in significantly higher communication costs and limited scalability. In contrast,MaTU incurs minimal overhead by transmitting a single “unified” task vector per client, supplemented by compact task modulators (a binary mask and scalar value per task).
@@
Scaling with more client tasks. We further analyze how MaTU handles an increasing number of tasks per client. Specifically, we conduct experiments on the $30$-task benchmark, varying each client’s task groups from $2$ to $30$, and report both (i) communication cost per round (in MB) and (ii) average normalized performance across tasks (vs individual task fine-tuning). As shown in Fig.5(a)MaTU introduces minimal communication overhead as client tasks scale. This efficiency arises from transmitting a single adapter per client, regardless of number of tasks, supplemented by lightweight task-specific modulators (a binary mask and a scalar value), unlike the multiple adapters required byMaT-FL and other baselines. Furthermore, in terms of model performance, Fig.5(b) shows that MaT-FL suffers from a sharp performance drop when clients are assigned more than $5$ tasks, highlighting its limitations in managing high task heterogeneity. In contrary, MaTU demonstrates to have a solid performance – even when clients train on over $15$ tasks – highlighting its ability to effectively disentangle tasks in the weight space through task-specific modulators.
@@
Highly Conflicting task groups. Here, we evaluate MaTU’s performance in scenarios with highly conflicting tasks within client’s task group. Using the $8$-task benchmark, we focus on the $3$ distinct task clusters identified in Fig.2. Specifically, we conduct experiments with $10$ clients, each assigned a fixed group of three tasks, designed to include no, $2$, or $3$ highly dissimilar tasks, referred to as no conflict, 2-conflict, and 3-conflict task groups, respectively. This setup enables us to assess how MaTU handles task conflicts on the clients and compare its performance against baselines. As shown in Fig. 6MaTU achieves consistently high performance with minimal drop (less than $6$%), unlike baselines such as FedPer, which struggles to train multi-task models even without conflicting task groups, and MaT-FL, which suffers from a more pronounced performance drop as task conflicts increase.
Figure 6: Impact of conflicting tasks on MaTU. Performance evaluation of ViT-B/32 on the $8$-task benchmark. We report normalized test set accuracy vs individual task fine-tuning. Federated parameters are set to $N$=$10$, $R$=$300$, $\xi$=$1$ and $E$=$1$.
@@ 5 Conclusions
In this work, we introduced MaTU, a MaT-FL approach designed to address task heterogeneity by relying on task vectors that creates a “unified” task vector – encapsulating all tasks – together with lightweight task modulators. By eliminating the need for clustering or client-specific weight storage, MaTU enables scalable and communication-efficient training in MaT-FL. Our evaluations across $30$ datasets demonstrated that MaTU not only achieves superior performance compared to state-of-the-art MaT-FL regimes but also delivers results comparable to per-task fine-tuning while significantly reducing communication overhead; thus highlighting MaTU’s potential under highly heterogeneous FL scenarios.

### v2

@@ Abstract
Federated Learning (FL) traditionally assumes homogeneous client tasks; however, in real-world scenarios, clients often specialize in diverse tasks, introducing task heterogeneity. To address this challenge, Many-Task FL (MaT-FL) has emerged, enabling clients to collaborate effectively despite task diversity. Existing MaT-FL approaches rely on client grouping or personalized layers, requiring the server to manage individual models and failing to account for clients handling multiple tasks. We propose MaTU, a MaT-FL approach that enables joint learning of task vectors across clients, eliminating the need for clustering or client-specific weight storage at the server. MaTU constructs a single “unified” task vector per client and leverages lightweight task-specific modulators — binary masks and scalar rescalers — to disentangle and adapt per-task behavior. This design ensures stateless server aggregation and naturally scales to large task sets with minimal communication overhead. Evaluated across 30 datasets, MaTU achieves superior performance over state-of-the-art MaT-FL approaches, with results comparable to per-task fine-tuning, while delivering significant communication savings.
@@ 1 Introduction
Recent studies Chen et al. (2023); Zhuang et al. (2023); Cai et al. (2023) have underscored the significance of task heterogeneity in FL. To address this diversity, traditional FL has evolved into Many-Task Federated Learning (MaT-FL) Cai et al. (2023) to enable clients to effectively collaborate, despite specializing in different tasks. It is worth to mention that while “many-task” has been interchangeably used as “multi-task” in conventional ML, early Multi-Task Federated Learning (MTFL) mainly addressed personalized FL, a subset of MaT-FL. MaT-FL is thus used to distinguish it from prior MTFL works Cai et al. (2023); Muhamed et al. (2024); Zhuang et al. (2023). Current MaT-FL approaches address task heterogeneity by focusing on client grouping or split FL. For instance, FedBone Chen et al. (2023) employs split FL whereas the server needs to sustain a unique model per client. Conversely, MaT-FL Cai et al. (2023) and MAS Zhuang et al. (2023) concentrate on dynamically grouping client models based on task similarity, aggregating models among clients that handle similar tasks. Nevertheless, these strategies require the server to manage individual models for each client or task and fail to address scenarios where clients hold more than one task ( e.g., a personal health tracker monitoring multiple biomarkers such as heart-rate, blood-pressure, blood glucose level, body temperature simultaneously ). Training a single model among tasks offers considerable benefits in federated settings. Training individual models for each task is resource-intensive, whereas a unified model not only conserves resources but also potentially enhances performance by leveraging learning from diverse auxiliary tasks.
@@
This paper addresses the challenge of adapting pretrained Foundation Models (FMs) to the MaT-FL setting, where clients hold an arbitrary (>$1$) number of tasks, and we aim to train a unified model across all tasks. We propose Many-Task FL via Unified Task Vectors (MaTU), a novel approach that eliminates the need for clustering or client-specific weight storage at the server, enabling natural knowledge transfer across tasks in FL. Inspired by recent advancements in model merging and Task Arithmetic (TA) Ilharco et al. (2023); Yadav et al. (2023a)MaTU introduces a task vector aggregation mechanism tailored for MaT-FL. Specifically, we determine task similarity directly based on the direction of clients’ task vectors and constructs a “unified” task vector that encapsulates all tasks. To handle task-specific requirements, we augment the “unified” vector with lightweight modulators (i.e., binary task-specific masks and scalers) that facilitate knowledge transfer between related tasks while promoting weight disentanglement for dissimilar ones, ensuring robust performance across heterogeneous tasks in FL. By transmitting only a single vector and lightweight modulators per client, MaTU avoids the per-task overhead of existing MaT-FL methods, enabling both communication efficiency and scalability to large task sets. Concisely, our main contributions are as follows:
@@ 2 Related Work
MaT-FL. Many-Task Federated Learning (MaT-FL) has emerged to tackle task heterogeneity in FL. From MTFL approaches, FedProx Li et al. (2020), introduced a proximity term to limit client updates from deviating too far from the global model in MTFL. Alternatively, FedBone Chen et al. (2023) used split FL to introduce task-specific personalized layers to handle multiple tasks in FL. Recent MaT-FL approaches primarily address task heterogeneity by focusing on client grouping Cai et al. (2023); Zhuang et al. (2023); Lu et al. (2023). MaT-FL Cai et al. (2023) and MAS Zhuang et al. (2023) dynamically group clients based on task similarity, enabling model aggregation among clients with similar tasks. FedHCA Lu et al. (2023) extends this by supporting a variable number of tasks across clients, creating a more flexible framework. More recently, NTKFedAvg Muhamed et al. (2024) introduced task arithmetic in MaT-FL, enabling clients to train task-specific adapters and leverage server-side adapter fusion for optimizing multiple tasks. To further improve task disentanglement during model aggregation, they applied Neural Tangent Kernel (NTK) linearization over the model prior to training. Nevertheless, group-based MaT-FL approaches require the server to manage multiple models for each task group, assume task relationships are known in advance, and introduce significant client-side overhead, making scalability a challenge as the number of tasks increases. Training a single model across tasks offers considerable benefits: it reduces resource consumption and can enhance performance by leveraging diverse auxiliary tasks. While NTKFedAvg attempts to address these challenges, it focuses on a single task per client, whereas in practice, clients may hold multiple tasks. MaTU bridges this gap by training a unified model across all clients and tasks, dynamically building task correlations to enable knowledge transfer, and using task vector aggregation based on task similarity, removing the need for multiple server-side models.
@@ 3.1 Preliminaries
For the set of tasks assigned to each clients, $\mathcal{T}_{n}$, we can now utilize the newly computed aggregated task-specific vector to derive their updated “unified” task vector, $\tau^{r+1}_{n}$, and subsequently compute the set of task-specific masks, $\mathcal{M}^{r+1}_{n}=\{(\tau^{i,r+1}\odot\tau^{r+1}_{n}>0)\}^{k_{n}}_{i=1}$, and scaling parameters $\lambda^{r+1}_{n}=\left\{\frac{\sum|\tau^{i,r+1}_{n}|}{\sum|m^{i,r+1}_{n}\odot\tau^{r+1}_{n}|}\right\}_{i=1}^{k_{n}}$. The client’s “unified” task vector, $\tau^{r+1}_{n}$, and task-specific modules, $\mathcal{M}^{r+1}_{n}$ and $\lambda^{r+1}_{n}$, are then transmitted to the client-side for subsequent training. To train task $t$, the client must modulate the “unified” vector as $\dot{\tau}^{r+1}_{t}=\lambda^{t,r+1}_{n}\cdot m^{t,r+1}_{n}\odot\tau^{r+1}_{n}$, and injects it to the pre-trained model (i.e., $\theta^{r+1}_{t}=\theta_{p}+\dot{\tau}^{r+1}_{t}$). In contrast to personalized FL or group-based MaT-FL methods that require storing task-specific heads or group models, our design remains stateless — the server only handles task-level aggregation and does not retain client-specific weights across rounds.
@@ 4.1 Results
$8$-task benchmark: Table 2 demonstrates that MaTU maintains strong performance in multiple-task settings, achieving an average accuracy of 79.47% – a modest 5% drop from single-task settings – while outperforming all baselines and effectively minimizing task interference via masking. Traditional FL methods, such as FedAvg and FedProx, exhibit significant performance deficits, with average accuracies well below their single-task counterparts, underscoring their limitations in addressing task heterogeneity. Notably, NTK-FedAvg surpasses these baselines, likely benefiting from improved task disentanglement through NTK-based linearization. In stark contrast, FedPer experiences a severe performance decline ($\approx$40%) when transitioning from single to multiple-task settings, reflecting its inability to manage intra-client task heterogeneity due to its reliance on personalization layers. This highlights that personalized FL approaches are not designed to address the broader challenges of task heterogeneity, underscoring that personalization represents a subset of the more general MaT-FL problem. Conversely, MaT-FL achieves higher performance by employing similarity-based grouping to mitigate task conflicts, reducing interactions among dissimilar clients and tasks, but still lags MaTU by 10%. We further analyze how MaT-FL and MaTU handle such conflicts in Fig.˜6(a).
Beyond test accuracy, MaTU offers substantial communication and storage efficiency in multiple-task settings. As shown by the $\mathrm{bpt}$ column in Table 2, competing methods transmit multiple adapters — one per task—leading to high communication costs and server-side storage of client- or group-specific weights. In contrast, MaTU transmits only a single unified task vector per client, along with lightweight modulators (a binary mask and a scalar per task), enabling stateless aggregation and scalable deployment without storing any client-specific parameters.
@@
Scaling clients tasks. We further analyze how MaTU handles an increasing number of tasks per client. Specifically, we conduct experiments on the $30$-task benchmark, varying each client’s task groups from $2$ to $30$, and report both (i) communication cost per round (in MB) and (ii) average normalized performance across tasks (vs individual task fine-tuning). As shown in Fig.5(a)MaTU introduces minimal communication overhead as client tasks scale. This efficiency arises from transmitting a single adapter per client, regardless of number of tasks, supplemented by lightweight task-specific modulators (a binary mask and a scalar value), unlike the multiple adapters required byMaT-FL and other baselines. Furthermore, in terms of model performance, Fig.5(b) shows that MaT-FL suffers from a sharp performance drop when clients are assigned more than $5$ tasks, highlighting its limitations in managing high task heterogeneity. In contrary, MaTU demonstrates to have a solid performance – even when clients train on over $15$ tasks – highlighting its ability to effectively disentangle tasks in the weight space through task-specific modulators.
@@
(a) Task conflict groups
(b) Cross-task aggregation
Figure 6: Performance evaluation of ViT-B/32 on the $8$-task benchmark under (a) task conflict groups and (b) cross-task aggregation variants. We report (a) normalized test set accuracy vs individual task fine-tuning and (b) average accuracy on test set. Federated parameters are set to $N$=$10$, $R$=$300$, $\xi$=$1$ and $E$=$1$.
Highly Conflicting task groups. Here, we evaluate MaTU’s performance in scenarios with highly conflicting tasks within client’s task group. Using the $8$-task benchmark, we focus on the $3$ distinct task clusters identified in Fig.2. Specifically, we conduct experiments with $10$ clients, each assigned a fixed group of three tasks, designed to include no, $2$, or $3$ highly dissimilar tasks, referred to as no conflict, 2-conflict, and 3-conflict task groups, respectively. This setup enables us to assess how MaTU handles task conflicts on the clients and compare its performance against baselines. As shown in Fig. 6(a)MaTU achieves consistently high performance with minimal drop (less than $6$%), unlike baselines such as FedPer, which struggles to train multi-task models even without conflicting task groups, and MaT-FL, which suffers from a more pronounced performance drop as task conflicts increase.
Cross-task aggregation. We also analyze the benefits of cross-task aggregation (Eq. 6) by comparing MaTU against two variants: one without cross-task aggregation and another using uniform cross-task averaging, where all task vectors are treated equally regardless of similarity. As shown in Fig.6(b)MaTU achieves superior performance in the absence of conflicts and effectively mitigates negative transfer under 2-conflict settings. In contrast, the uniform variant degrades under task heterogeneity, while removing cross-task aggregation entirely limits performance.
@@ 5 Conclusions
In this work, we introduced MaTU, a MaT-FL approach designed to address task heterogeneity by relying on task vectors that creates a “unified” task vector – encapsulating all tasks – together with lightweight task modulators. By eliminating the need for clustering or client-specific weight storage, MaTU  enables scalable, low-overhead MaT-FL even with large numbers of tasks. Our evaluations across $30$ datasets demonstrated that MaTU not only achieves superior performance compared to state-of-the-art MaT-FL regimes but also delivers results comparable to per-task fine-tuning while significantly reducing communication overhead; thus highlighting MaTU’s potential under highly heterogeneous FL cases.
Acknowledgments
This work was funded by the DAIS project, which has received funding from ECSEL Joint Undertaking under grant agreement No 101007273, and the Bio-Curity project, which has received funding from Eureka Cluster Xecs under grant agreement No 2022016.